OpenStudy (firejay5):

Please help me with Math Problem: x = y over y + 4; solve for y

OpenStudy (pfenn1):

What do you need to do with this equation?

OpenStudy (firejay5):

solve for y

OpenStudy (pfenn1):

So you have to rearrange the equation to solve for y in terms of x?

OpenStudy (firejay5):

solve problem for y, cause of 2 y's

OpenStudy (firejay5):

there are 2 y's, one one each side I suppose

OpenStudy (pfenn1):

Okay your equation is:\[x=\frac{y}{y+4}\]If you multiple both sides by (y+4) you get\[x(y+4)=\frac{y\cancel{(y+4)}}{\cancel{y+4}}\]or\[xy+4x=y\]Now subtract xy from both sides to get\[\cancel{xy}+4x-\cancel{xy}=y-xy\]or\[4x=(1-x)y\]

OpenStudy (firejay5):

thank you so much! :D best answer ever

OpenStudy (pfenn1):

you are welcome.

OpenStudy (pfenn1):

You can rearrange that last equation to show\[ y=\frac{4x}{1-x}\]

OpenStudy (firejay5):

so what's the official answer to it

OpenStudy (pfenn1):

This last one....it is the same as the one I gave before. I just divided both sides by 1-x.\[\frac{4x}{1-x}=\frac{y\cancel{(1-x)}}{\cancel{1-x}}\]

OpenStudy (firejay5):

where did you get 1

OpenStudy (firejay5):

is my question

OpenStudy (pfenn1):

Oh. From this part of the calculation\[\cancel{xy}+4x-\cancel{xy}=y-xy =>4x=y-xy\]\[y-xy=y(1-x) \]Does that make sense?